A cracker rocket is ejecting 0.05 kg of gases per sec at a velocity of 400m/s.The accelerating force on the rocket is


a.20 dynes


b.20 N


c.20 kg-wt.


d.none of the above.

Grade:12th Pass

2 Answers

Rahul Kumar
12 years ago

Momentum of the ejected gases= mass*velocity

=0.05*400=20kgm/s

 

Hence accelerating force of the rocket is 20 N

Akshil
5 years ago
Momentum of the ejection gases=mass*velocity
           =0.05*400
           =20kgm/s
           =20N
 Hence the acceleration force of the rocket is 20 N